Two students are going to the store to buy school supplies. One of the students buys 2 packs of pencils and 3 packs of pens for $8.25. Her friend purchases 5 packs of pencils and 2 packs of pens for $11. What is the cost of 1 pack of pencils and 1 pack of pens?

Answer by Boreal(15235)

pencils=x
pens=y
2x+3y=8.25
5x+2y=11
multiply the top by 2 and the bottom by -3 and add, eliminating y
4x+6y=16.50
-15x-6y=-33
-11x=-16.50
x=$1.50 PENCIL
substitute into first
3+3y=8.25
3y=5.25
y=$1.75 PEN
Check in second
7.50+3.50=$11 checks.